### v1

# AMNS: Attention-Weighted Selective Mask and Noise Label Suppression for Text-to-Image Person Retrieval thanks: †: Corresponding author

Abstract
Text-to-image person retrieval aims to retrieve images of person given textual descriptions, and most methods implicitly assume that the training image-text pairs are correctly aligned, but in practice, under-correlated and false-correlated problems arise for image-text pairs due to poor image quality and mislabeling. Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy pairings between image lexical elements and text descriptions. To solve these two problems, we propose a new noise label suppression method and alleviate the problem generated by random mask through an attention-weighted selective mask strategy. In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focus loss to further improve the model’s recognition ability on difficult samples. On the other hand, Attention-Weighted Selective Mask processes the raw image through the EMA version of the image encoder, retaining some of the tokens with strong semantic associations with the corresponding text descriptions in order to extract better features. Numerous experiments validate the effectiveness of our approach in terms of dealing with noisy problems. The code will be available soon at https://github.com/RunQing715/AMNS.git.
Index Terms:
Text-to-image person retrieval, cross-modal retrieval, noise suppression, image mask
I Introduction
Text-to-image person retrieval utilizes textual descriptions to aid in the retrieval of person images [1, 2, 3]. Although many effective retrieval methods have been proposed in the field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input training pairs are correctly aligned. However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, making it difficult for the idealized assumptions to hold. In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noise matching between image patches and text description, and ultimately lead to the degradation of retrieval performance. Therefore, effective noise suppression methods and mask strategies for images are crucial for improving the performance of model retrieval.
In particular, some images are mislabeled, or the image quality is too low, but are still treated as positive examples for cross-modal learning, resulting in the model learning the wrong correspondences, a phenomenon known as Noisy Correspondence (NC). To solve this problem, there are two main approaches: sample selection[14, 15, 16, 17] and robust loss functions[18, 19, 20]. The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carries the risk of mistakenly deleting valid data. Comparatively, the latter approach aims to develop noise-tolerance loss functions to improve the robustness of model training against NC.
Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which effectively suppresses the negative impact of noisy labels by taking into account the bidirectional relationship between the predictive distribution and the true distribution, and is complemented by the introduction of a Weight Adjustment Focus (WAF) loss to enhance the model’s ability to handle difficult samples. In addition, to address the problem that random mask may destroy the semantic association between graphs and texts, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy pairings. The main contributions and innovations of this paper are as follows:
  • We investigate two new and prevalent problems in text-to-image person retrieval tasks, namely, the Noisy Correspondence (NC) and the noisy pairing problem, which may mislead the model to learn incorrect semantic information. To the best of our knowledge, this paper may be the first work to explore the noise pairing problem caused by the random masking strategy in person retrieval tasks.
  • We present AMNS: attention-weighted selective mask and noise label suppression framework, by introducing Attention-Weighted Selective Mask (AWM) strategy to alleviate the noise pairing problem between graphs and texts caused by random masks, and by proposing Bidirectional Similarity Distribution Matching (BSDM) loss is proposed to effectively suppress the negative impact caused by noisy labels, and Weight Adjustment Focal (WAF) loss is proposed to be combined with BSDM loss, which further improves the model’s recognition ability on difficult samples.
  • Experiments show that using AWM strategy outperforms random masking strategy, and that the integration of the proposed method for suppressing noisy labels into other researchers’ methods achieves an overall improvement in performance. Meanwhile, both the mAP index and mINP index of AMNS are almost the best among the comparison methods, which fully proves the superiority of our proposed method and can be better used in real-life retrieval systems.
Figure 1: An overview of our AMNS.
II METHOD
The overall diagram of our proposed framework is shown in Fig. 1, which consists of an image encoder ${\mathbf{E_{i}}}$, a text encoder ${\mathbf{E_{t}}}$ and an EMA-based image encoder $\overline{\mathbf{E_{i}}}$. Given an image, crop it and input it to $\overline{\mathbf{E_{i}}}$, a mask image is obtained using an AWM strategy to highlight image regions relevant to the textual description and fed into ${\mathbf{E_{i}}}$ together with the raw image, enabling ${\mathbf{E_{i}}}$ to extract a more accurate representation of the features. The model makes $\overline{\mathbf{E_{i}}}$ filter out features with lower weights by self-supervised learning, and combines with the proposed BSDM, which considers the inverse KL scatter, to inhibit the noisy labelling problem by inversely constraining the true label distribution through the predictive distribution. Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of difficult samples by adjusting the weight coefficients. In addition, we adopt ID loss to aggregate feature representations of the same identity, which further improves the model retrieval performance.
Figure 2: BSDM diagram. It demonstrates the core idea of implementing noisy label suppression. This approach constrains the true label distribution through the predicted distribution, preventing the model from becoming overconfident and mitigating the impact of noisy labels.
II-A Attention-Weighted Selective Mask
[…]
$\mathit{A}_{\mathit{P}}=\frac{1}{HL}\sum_{l=1}^{L}\sum_{h=1}^{H}\text{Softmax}\left(\frac{\mathbf{f}_{lh}^{q}(\mathrm{CLS})\cdot\mathbf{f}_{lh}^{k}(P)}{\sqrt{d_{k}}}\right)$ (1)
[…]
For a complete image, we introduce an Exponential Moving Average (EMA) version of the image encoder to generate the attention weights. At the same time, we use the Multiple Views strategy[22], where we take two views and continue to introduce two self-supervised learning methods (SimCLR loss[23] and BYOL loss[24]) to assist in this mask task. Overall, the total loss of the AWM strategy is:
$L_{AWM}=L_{ITC}+L_{SimCLR}+L_{BYOL}$ (2)
II-B Noise Label Suppression Methods
SDM loss[6] effectively differentiates matching from non-matching pairs by aligning the cosine similarity distribution $\mathbf{p}$ of image-text pairs with the ideal label distribution $\mathbf{q}$. However, due to noisy correspondences in the dataset, SDM forcibly fitting all positive sample pairs of a dataset causes the model to overfit data containing noise.
Therefore, we propose Bidirectional Similarity Distribution Matching (BSDM), the core idea of which is shown in Fig. 2. Building upon the forward KL divergence, BSDM incorporates the inverse KL divergence, so that the distribution $\mathbf{p}$ can also exert constraints on the distribution $\mathbf{q}$. This bidirectional constraint helps alleviate Noisy Correspondences (NC) by allowing the model some flexibility in fitting the data, rather than blindly striving for perfect alignment with all positive pairs.
0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.873.073.073.073.073.573.573.573.574.074.074.074.074.574.574.574.575.075.075.075.0Mask RatioRank-1AMNSAWMAMNSRandom
Figure 3: Variation of Rank-1 performance with different mask ratios for $m$ and $\tau$.
First, given $N$ image-text pairs, $f_{i}^{v}$ and $f_{j}^{t}$ denote image global features and text global features. We construct a set of image-text representation pairs as $\{(f_{i}^{\nu},f_{j}^{t}),y_{i,j}\}_{j=1}^{N}$, where $y_{i,j}=1$ represents matched pairs from the same identity, and $y_{i,j}=0$ represents unmatched pairs, $sim(\mathbf{u},\mathbf{v})=\mathbf{u}^{\mathsf{T}}\mathbf{v}/\|\mathbf{u}\|\|\mathbf{v}\|$ denotes the cosine similarity between u, v, and $\tau$ is the temperature hyperparameter. Then the probability of matching pairs can be calculated using the following softmax function.
$p_{i,j}=\frac{\exp{(sim(f_{i}^{v},f_{j}^{t})/\tau)}}{\sum_{k=1}^{N}\exp{(sim(f_{i}^{v},f_{k}^{t})/\tau)}}$ (3)
The image-to-text $\mathcal{L}_{i2t}$ loss in BSDM is computed by bidirectional KL scattering:
$\mathcal{L}_{i2t}$ $=KL(\mathbf{p_{i}}\parallel\mathbf{q_{i}})+KL(\mathbf{q_{i}}\parallel\mathbf{p_{i}})$ (4)
$=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\left(p_{i,j}\log\left(\frac{p_{i,j}}{q_{i,j}+\epsilon}\right)+q_{i,j}\log\left(\frac{q_{i,j}+\epsilon}{p_{i,j}}\right)\right)$
where $\mathbf{\epsilon}$ is used to avoid very small numbers in numerical problems. $q_{i,j}=y_{i,j}/\sum_{k=1}^{N}y_{i,k}$ is the true matching probability. $KL(\mathbf{p_{i}}\parallel\mathbf{q_{i}})$ is the forward KL scatter, and the goal is to make $\mathbf{p_{i}}$ as close to $\mathbf{q_{i}}$ as possible, this helps the model to learn the correct data representation. $KL(\mathbf{q_{i}}\parallel\mathbf{p_{i}})$ is the inverse KL scatter. With the bidirectional KL scattering, the model is made to be pushed not only to adapt to the real labels that may contain noise, but also to allow the real labels to adapt to the model’s predictions, thus mitigating the effects of NC.
Correspondingly, the loss from text to image $\mathcal{L}_{t2i}$ can be exchanged for $f^{v}$ and $f^{t}$ in Eq. (3) (4), BSDM loss is calculated by:
$\mathcal{L}_{BSDM}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}$ (5)
TABLE I: Comparison of the performance of the three losses SDM, TAL, and BSDM in the IRRA[6], RDE[37], and AMNS methods on the CUHK-PEDES dataset, where "Best" denotes the selection of the best checkpoints to be tested on the validation set, and the best results are shown in bold. The RDE method continues its performance comparison using two synthetic noise rates, i.e., 20% and 50% to simulate the performance under the real field where the image-text pairs are not well aligned.
  Loss CUHK-PEDES
Noise Methods SDM TAL BSDM Rank-1 Rank-5 Rank-10 mAP mINP
IRRASDM Best 73.38 89.93 93.71 66.13 50.24
0% IRRATAL Best 73.20 89.02 93.34 66.05 50.49
IRRABSDM Best 73.51 89.38 93.75 66.27 50.71
RDESDM Best 75.26 89.91 94.02 67.37 51.40
0% RDETAL Best 75.94 90.14 94.12 67.56 51.44
RDEBSDM Best 76.06 90.38 94.36 68.04 51.99
RDESDM Best 74.27 89.49 93.62 66.34 50.10
20% RDETAL Best 74.46 89.42 93.63 66.13 49.66
RDEBSDM Best 74.95 90.03 94.02 66.88 50.84
RDESDM Best 69.33 86.99 91.68 61.99 45.34
50% RDETAL Best 71.33 87.41 91.81 63.50 47.36
RDEBSDM Best 71.33 87.30 92.20 63.75 47.51
AMNSSDM Best 73.73 89.18 93.55 67.35 52.88
0% AMNSTAL Best 73.39 88.97 93.13 67.17 52.82
AMNSBSDM Best 74.25 89.54 93.83 67.67 53.16
BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for difficult samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples). Therefore, we consider the introduction of focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ we are able to adjust the model’s attention to the misclassified samples. We refer to this loss as Weight Adjustment Focal Loss (WAF). The image-to-text WAF loss is as follows:
$\mathcal{L}_{W\!AF-i2t}=\begin{cases}-\alpha(1-p_{i,j})^{\gamma}\log p_{i,j},&y_{i,j}=1\\ -\beta p_{i,j}^{\gamma}\log(1-p_{i,j}),&y_{i,j}=0\end{cases}$ (6)
Correspondingly, the total WAF loss is as follows:
$\mathcal{L}_{W\!AF}=\mathcal{L}_{W\!AF-i2t}+\mathcal{L}_{W\!AF-t2i}$ (7)
By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of difficult samples, thus achieving better performance in noisy data environments.
Optimization.We also employ the commonly used ID loss[36], whose main purpose is to make feature representations of the same identity clustered together in the feature space, while feature representations of different identities are separated. This helps the model to better distinguish between different entities, thus improving the performance of model retrieval. The overall optimization objective of our training is defined as:
$\mathrm{L}=\mathrm{L}_{AWM}+\mathrm{L}_{BSDM}+\mathrm{L}_{W\!AF}+\mathrm{L}_{id}$ (8)
TABLE II: Performance comparison on three datasets. The best and second-best results are in bold and underline, respectively.
  CUHK-PEDES ICFG-PEDES RSTPReid
Methods Ref. Rank-1 Rank-5 Rank-10 mAP mINP Rank-1 Rank-5 Rank-10 mAP mINP Rank-1 Rank-5 Rank-10 mAP mINP
IVT[30] ECCVW’22 65.59 83.11 89.21 - - 56.04 73.60 80.22 - - 46.70 70.00 78.80 - -
CFine[32] TIP’23 69.57 85.93 91.15 - - 60.83 76.55 82.42 - - 50.55 72.50 81.60 - -
PBSL[33] ACMMM’23 65.32 83.81 89.26 - - 57.84 75.46 82.15 - - 47.80 71.40 79.90 - -
BEAT[34] ACMMM’23 65.61 83.45 89.54 - - 58.25 75.92 81.96 - - 48.10 73.10 81.30 - -
LCR2S[19] ACMMM’23 67.36 84.19 89.62 59.24 - 57.93 76.08 82.40 38.21 - 54.95 76.65 84.70 40.92 -
IRRA[6] CVPR’23 73.38 89.93 93.71 66.13 50.24 63.46 80.25 85.82 38.06 7.93 60.20 81.30 88.20 47.17 25.28
RDE[37] CVPR’24 75.94 90.14 94.12 67.56 51.44 67.68 82.47 87.36 40.06 7.87 65.35 83.95 89.90 50.88 28.08
CSKT[8] ICASSP’24 69.70 86.92 91.80 62.74 - 58.90 77.31 83.56 33.87 - 57.75 81.30 88.35 46.43 -
Our AMNS - 74.25 89.54 93.83 67.67 53.16 64.05 79.90 85.90 41.27 10.36 60.50 81.30 87.65 47.20 25.38
III Experimental section
III-A Datasets and Settings
Datasets: In our experiments, we evaluated the proposed methods on CUHK-PEDES[1], ICFG-PEDES[26], and RSTPReid[27] datasets.
Evaluation Protocols: For all experiments, we use the popular Rank-K metric (K=1,5,10) to measure retrieval performance. We also use mean Average Precision (mAP) and mean Inverse Negative Penalty (mINP[28]) as auxiliary retrieval metrics to further evaluate the performance.
Implementation Details: In our experiments, we used the same version of CLIP-ViTB/16 as IRRA[6]. Image augmentation includes color jitter, random horizontal flipping, random crop with padding, random grayscale, random erasing, and gaussian blur. Text augmentation involves random masking, replacement, and removal of word tokens. The model training parameters follow IRRA, the temperature parameter $\tau$ of BSDM is set to 0.02, $\gamma$ of WAF is set to 2, and the values of $\alpha$ and $\beta$ are set to 0.1 and 0.05, respectively. $\overline{\mathbf{E_{i}}}$ generates an attention mask with a momentum that starts from 0.996 and gradually increases to 1 with training, and the experiments show that it works best when the mask rate is 0.5. Specific experimental results are detailed in Section III-B.
III-B Experimental results
In Fig. 3, the experimental results on AMNS show that the effect of attention-weighted selective mask based on attention is generally better than that of random mask, and the effect of attention-weighted selective mask reaches the optimal effect when the mask ratio = 0.5. Therefore, in the subsequent experiments, for the AWM strategy, we all set the mask ratio to 0.5 for the experiments.
As shown in Table I, the SDM loss is the similarity distribution matching loss proposed by IRRA[6], which is introduced in Section II-B. TAL loss is a novel triplet alignment loss proposed by RDE[37], which effectively solves the problem of poor local minima and even model collapse under NC in the early stage of training and achieves better performance.The BSDM is the bidirectional similarity distribution matching loss proposed in this paper. Table I shows that the use of BSDM loss in different methods can bring significant improvement in all aspects of performance, and the best performance is still achieved under the 20% and 50% noise rates, which fully indicates that the BSDM loss proposed by us is superior to the widely used TAL loss and SDM loss, and has better stability and robustness for NC.
Table II shows that our method achieves good results in Rank metrics, and the mAP index and mINP index achieve the best results in the first and second datasets. Higher mINP metrics mean that the correctly matched images are indexed at a more advanced position, which fully validates that our method is better at finding all the correct matches and can be better used in real-life retrieval systems.
TABLE III: Ablation study on each component of AMNS on CUHK-PEDES.
  Components CUHK-PEDES
No. Methods BSDM WAF Lid Rank-1 Rank-5 Rank-10 mAP mINP
0 Baseline 69.12 87.02 92.02 62.17 46.08
1 +BSDM 73.34 89.23 93.02 65.25 48.76
2 +WAF 71.74 88.09 92.56 64.94 49.35
3 +Lid 66.57 85.62 91.00 61.12 45.82
4 +WAF+Lid 69.83 86.11 91.55 65.28 52.07
5 +BSDM+WAF 73.60 88.81 93.11 65.56 49.21
6 +BSDM+Lid 74.16 89.02 93.52 67.37 52.57
7 AMNS 74.25 89.54 93.83 67.67 53.16
III-C Ablation experiments
In this section, we fine-tune the CLIP-ViT-B/16 model with AWM loss as a baseline and perform ablation studies on the CUHK-PEDES dataset.
Effectiveness of BSDM: As can be seen from No.1 in Table III, just by adding BSDM to Baseline, the accuracies of Rank-1, Rank-5, and Rank-10 are improved by 4.22%, 2.21%, and 1.00%, respectively, and the values of mAP and mINP are improved by 3.08% and 2.68%, respectively, which is a good illustration of the BSDM’s Person Retrieval task in terms of its Effectiveness.
Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize difficult samples when it is complemented with BSDM.
IV Conclusion
In this paper, we explore Noise Correspondence (NC) and the noise pairing problem caused by random masking strategy. We introduce Attention-Weighted Selective Mask to alleviate the noise matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focus loss, effectively deal with the NC, and achieve excellent performance. We have carried out extensive experiments, which fully prove the superiority of the proposed method in solving the above problems, and prove that our method can find all the correct matches better.

### v2

@@
 Text-to-image person retrieval aims to retrieve images of person given textual descriptions, and most methods implicitly assume that the training image-text pairs are correctly aligned, but in practice, under-correlated and false-correlated problems arise for image-text pairs due to poor image quality and mislabeling.
-Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy pairings between image lexical elements and text descriptions.
+Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy matching between image lexical elements and text descriptions.
 To solve these two problems, we propose a new noise label suppression method and alleviate the problem generated by random mask through an attention-weighted selective mask strategy.
-In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focus loss to further improve the model’s recognition ability on difficult samples.
+In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focal loss to further improve the model’s recognition ability on hard samples.
 On the other hand, Attention-Weighted Selective Mask processes the raw image through the EMA version of the image encoder, retaining some of the tokens with strong semantic associations with the corresponding text descriptions in order to extract better features.
@@
 Although many effective retrieval methods have been proposed in the field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input training pairs are correctly aligned.
-However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, making it difficult for the idealized assumptions to hold.
-In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noise matching between image patches and text description, and ultimately lead to the degradation of retrieval performance.
-Therefore, effective noise suppression methods and mask strategies for images are crucial for improving the performance of model retrieval.
+However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, etc., making it difficult for the idealized assumptions to hold.
+In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noisy matching between image patches and text description, and ultimately lead to the degradation of retrieval performance.
+Therefore, effective noise suppression methods and image masking strategy are crucial for improving the performance of model retrieval.
 In particular, some images are mislabeled, or the image quality is too low, but are still treated as positive examples for cross-modal learning, resulting in the model learning the wrong correspondences, a phenomenon known as Noisy Correspondence (NC).
 To solve this problem, there are two main approaches: sample selection[14, 15, 16, 17] and robust loss functions[18, 19, 20].
-The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carries the risk of mistakenly deleting valid data.
+The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carry the risk of mistakenly deleting valid data.
 Comparatively, the latter approach aims to develop noise-tolerance loss functions to improve the robustness of model training against NC.
-Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which effectively suppresses the negative impact of noisy labels by taking into account the bidirectional relationship between the predictive distribution and the true distribution, and is complemented by the introduction of a Weight Adjustment Focus (WAF) loss to enhance the model’s ability to handle difficult samples.
-In addition, to address the problem that random mask may destroy the semantic association between graphs and texts, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy pairings.
+Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which introduces the inverse KL scattering while retaining the forward KL scattering, so that the model does not blindly pursue a perfect match with all the pairs of positive samples when fitting the data, but rather maintains a certain degree of flexibility, thus effectively suppressing the negative impact of noise labeling.
+And it is complemented by the introduction of a Weight Adjustment Focal Loss (WAF) based to enhance the model’s ability to handle hard samples.
+In addition, to address the problem that random mask may destroy the semantic association between image patches and text description, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy matching.
 The main contributions and innovations of this paper are as follows:
-  • We investigate two new and prevalent problems in text-to-image person retrieval tasks, namely, the Noisy Correspondence (NC) and the noisy pairing problem, which may mislead the model to learn incorrect semantic information.
-To the best of our knowledge, this paper may be the first work to explore the noise pairing problem caused by the random masking strategy in person retrieval tasks.
-  • We present AMNS: attention-weighted selective mask and noise label suppression framework, by introducing Attention-Weighted Selective Mask (AWM) strategy to alleviate the noise pairing problem between graphs and texts caused by random masks, and by proposing Bidirectional Similarity Distribution Matching (BSDM) loss is proposed to effectively suppress the negative impact caused by noisy labels, and Weight Adjustment Focal (WAF) loss is proposed to be combined with BSDM loss, which further improves the model’s recognition ability on difficult samples.
-  • Experiments show that using AWM strategy outperforms random masking strategy, and that the integration of the proposed method for suppressing noisy labels into other researchers’ methods achieves an overall improvement in performance.
-Meanwhile, both the mAP index and mINP index of AMNS are almost the best among the comparison methods, which fully proves the superiority of our proposed method and can be better used in real-life retrieval systems.
+  • We investigate two problems in text-to-image person retrieval tasks, namely the Noisy Correspondence (NC) and the noisy matching problem caused by random mask strategy.
+To our knowledge, this paper may be the first to explore the noisy matching problem caused by random mask strategies in pedestrian retrieval tasks.
+  • We present AMNS: attention-weighted selective mask and noise label suppression framework and introduce an Attention-Weighted Selective Mask (AWM) strategy to alleviate the noisy matching problem caused by random masks.
+It is proposed that Bidirectional Similarity Distribution Matching (BSDM) can effectively suppress the influence of noise labels, and combined with Weight Adjustment Focal Loss (WAF), The recognition ability of the model on difficult samples is improved.
+  • Experiments show that AWM strategy is superior to random masking strategy.
+In addition, the noise suppression method proposed in this paper is applied to other models, and the performance is significantly improved.
+Among the AMNS methods, mAP index and mINP index are almost the best.
 Figure 1: An overview of our AMNS.
@@
 The model makes $\overline{\mathbf{E_{i}}}$ filter out features with lower weights by self-supervised learning, and combines with the proposed BSDM, which considers the inverse KL scatter, to inhibit the noisy labelling problem by inversely constraining the true label distribution through the predictive distribution.
-Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of difficult samples by adjusting the weight coefficients.
+Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of hard samples by adjusting the weight coefficients.
 In addition, we adopt ID loss to aggregate feature representations of the same identity, which further improves the model retrieval performance.
@@
 AMNSBSDM Best 74.25 89.54 93.83 67.67 53.16
-BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for difficult samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples).
+BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for hard samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples).
 Therefore, we consider the introduction of focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ we are able to adjust the model’s attention to the misclassified samples.
@@
 $\mathcal{L}_{W\!AF}=\mathcal{L}_{W\!AF-i2t}+\mathcal{L}_{W\!AF-t2i}$ (7)
-By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of difficult samples, thus achieving better performance in noisy data environments.
+By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of hard samples, thus achieving better performance in noisy data environments.
 Optimization.We also employ the commonly used ID loss[36], whose main purpose is to make feature representations of the same identity clustered together in the feature space, while feature representations of different identities are separated.
@@
 Effectiveness of BSDM: As can be seen from No.1 in Table III, just by adding BSDM to Baseline, the accuracies of Rank-1, Rank-5, and Rank-10 are improved by 4.22%, 2.21%, and 1.00%, respectively, and the values of mAP and mINP are improved by 3.08% and 2.68%, respectively, which is a good illustration of the BSDM’s Person Retrieval task in terms of its Effectiveness.
-Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize difficult samples when it is complemented with BSDM.
+Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize hard samples when it is complemented with BSDM.
 IV Conclusion
-In this paper, we explore Noise Correspondence (NC) and the noise pairing problem caused by random masking strategy.
-We introduce Attention-Weighted Selective Mask to alleviate the noise matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focus loss, effectively deal with the NC, and achieve excellent performance.
+In this paper, we explore Noise Correspondence (NC) and the noisy matching problem caused by random masking strategy.
+We introduce Attention-Weighted Selective Mask to alleviate the noisy matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focal loss, effectively deal with the NC, and achieve excellent performance.
 We have carried out extensive experiments, which fully prove the superiority of the proposed method in solving the above problems, and prove that our method can find all the correct matches better.

### v3

@@
 Abstract
-Text-to-image person retrieval aims to retrieve images of person given textual descriptions, and most methods implicitly assume that the training image-text pairs are correctly aligned, but in practice, under-correlated and false-correlated problems arise for image-text pairs due to poor image quality and mislabeling.
-Meanwhile, the random masking augmentation strategy may incorrectly discard semantic content resulting in the problem of generating noisy matching between image lexical elements and text descriptions.
-To solve these two problems, we propose a new noise label suppression method and alleviate the problem generated by random mask through an attention-weighted selective mask strategy.
-In the proposed noise label suppression method, the effect of noise labels is suppressed by preventing the model from being overconfident by considering the inverse KL scatter loss, which is combined with the weight adjustment focal loss to further improve the model’s recognition ability on hard samples.
-On the other hand, Attention-Weighted Selective Mask processes the raw image through the EMA version of the image encoder, retaining some of the tokens with strong semantic associations with the corresponding text descriptions in order to extract better features.
-Numerous experiments validate the effectiveness of our approach in terms of dealing with noisy problems.
-The code will be available soon at https://github.com/RunQing715/AMNS.git.
+Most existing text-to-image person retrieval methods usually assume that the training image-text pairs are perfectly aligned; however, the noisy correspondence(NC) issue (i.e., incorrect or unreliable alignment) exists due to poor image quality and labeling errors.
+Additionally, random masking augmentation may inadvertently discard critical semantic content, introducing noisy matches between images and text descriptions.
+To address the above two challenges, we propose a noise label suppression method to mitigate NC and an Attention-Weighted Selective Mask (AWM) strategy to resolve the issues caused by random masking.
+Specifically, the Bidirectional Similarity Distribution Matching (BSDM) loss enables the model to effectively learn from positive pairs while preventing it from over-relying on them, thereby mitigating the risk of overfitting to noisy labels.
+In conjunction with this, Weight Adjustment Focal (WAF) loss improves the model’s ability to handle hard samples.
+Furthermore, AWM processes raw images through an EMA version of the image encoder, selectively retaining tokens with strong semantic connections to the text, enabling better feature extraction.
+Extensive experiments demonstrate the effectiveness of our approach in addressing noise-related issues and improving retrieval performance.
 Index Terms:
@@
 Text-to-image person retrieval utilizes textual descriptions to aid in the retrieval of person images [1, 2, 3].
-Although many effective retrieval methods have been proposed in the field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input training pairs are correctly aligned.
+Although many effective retrieval methods have been proposed in this field[4, 5, 6, 7, 8, 9, 10], almost all of them implicitly assume that all input image-text training pairs are correctly aligned.
 However, in the real world, noise is prevalent due to factors such as environmental disturbances, imprecise linguistic descriptions, mislabeling, or missing information, etc., making it difficult for the idealized assumptions to hold.
-In addition, the random masking strategy[11, 12, 13] in the process of image enhancement may incorrectly discard semantic content, resulting in noisy matching between image patches and text description, and ultimately lead to the degradation of retrieval performance.
-Therefore, effective noise suppression methods and image masking strategy are crucial for improving the performance of model retrieval.
-In particular, some images are mislabeled, or the image quality is too low, but are still treated as positive examples for cross-modal learning, resulting in the model learning the wrong correspondences, a phenomenon known as Noisy Correspondence (NC).
-To solve this problem, there are two main approaches: sample selection[14, 15, 16, 17] and robust loss functions[18, 19, 20].
+In addition, the random masking strategy[11, 12, 13] in the image enhancement process may incorrectly discard semantic content, leading to noisy matching between image patches and text descriptions, ultimately degrading retrieval performance.
+Therefore, effective noise suppression methods and image masking strategy are crucial for improving the model retrieval performance.
+Among these challenges, the issue of noisy image-text pairs is particularly prominent.
+Mislabeled or low-quality images, treated as positive examples, mislead the model into learning incorrect associations, this phenomenon known as Noisy Correspondence (NC).
+In order to solve the NC problem, there are two main approaches: sample selection[14, 15, 16, 17] and adopting robust loss functions[18, 19, 20].
 The former usually utilizes the memory effect of the deep neural networks[21] to progressively differentiate and filter noisy data, but this may require additional computational resources and carry the risk of mistakenly deleting valid data.
-Comparatively, the latter approach aims to develop noise-tolerance loss functions to improve the robustness of model training against NC.
-Therefore, this paper proposes a new cross-modal matching loss named image-text Bidirectional Similarity Distribution Matching (BSDM) loss, which introduces the inverse KL scattering while retaining the forward KL scattering, so that the model does not blindly pursue a perfect match with all the pairs of positive samples when fitting the data, but rather maintains a certain degree of flexibility, thus effectively suppressing the negative impact of noise labeling.
-And it is complemented by the introduction of a Weight Adjustment Focal Loss (WAF) based to enhance the model’s ability to handle hard samples.
-In addition, to address the problem that random mask may destroy the semantic association between image patches and text description, this paper introduces an Attention-Weighted Selective Mask (AWM) strategy, which aims to optimize the feature extraction process and improve the retrieval performance by accurately retaining the image regions that are strongly correlated with the textual descriptions and reducing the noisy matching.
+Comparatively, robust loss functions provide a more direct and efficient solution to address NC.
+Therefore, we focus on designing a noise-tolerant loss function to improve the model’s robustness.
+Specifically, blindly pursuing a perfect match with all positive pairs can cause overfitting on noisy labels.
+To address this, we propose a novel cross-modal matching loss called Bidirectional Similarity Distribution Matching (BSDM) loss.
+By integrating the cosine similarity distribution of image-text pairs into both forward and inverse KL divergence, BSDM helps the model associate representations across modalities more flexibly, reducing sensitivity to noise.
+Additionally, we introduce a Weight Adjustment Focal (WAF) loss to further enhance the model’s ability to handle hard samples, complementing BSDM in mitigating noisy labels’ effects.
+In addition, to address the problem that random mask may destroy the semantic association between image patches and text description, we introduce an Attention-Weighted Selective Mask (AWM) strategy.
+By computing attention weights for image regions, lower-weight regions are filtered out, allowing the model to focus on extracting more meaningful features.
+The AWM strategy aims to enhance feature extraction and retrieval performance, preserving image regions closely related to the text, while reducing noisy matches.
 The main contributions and innovations of this paper are as follows:
-  • We investigate two problems in text-to-image person retrieval tasks, namely the Noisy Correspondence (NC) and the noisy matching problem caused by random mask strategy.
-To our knowledge, this paper may be the first to explore the noisy matching problem caused by random mask strategies in pedestrian retrieval tasks.
-  • We present AMNS: attention-weighted selective mask and noise label suppression framework and introduce an Attention-Weighted Selective Mask (AWM) strategy to alleviate the noisy matching problem caused by random masks.
-It is proposed that Bidirectional Similarity Distribution Matching (BSDM) can effectively suppress the influence of noise labels, and combined with Weight Adjustment Focal Loss (WAF), The recognition ability of the model on difficult samples is improved.
+  • We investigate two problems in text-to-image person retrieval tasks, namely the Noisy Correspondence (NC) and the noisy matching problem caused by random masking strategy.
+To our knowledge, this paper may be the first to explore the noisy matching problem caused by random masking strategy in pedestrian retrieval tasks.
+  • We propose the AMNS framework.
+Specifically, Attention-Weighted Selection Mask (AWM) strategy alleviates noise matching caused by random masks.
+Bidirectional Similarity Distribution Matching (BSDM) loss enables effective learning from positive pairs while avoiding over-relying, mitigating overfitting to noisy labels.
+Additionally, Weight Adjustment Focal (WAF) loss improves the model’s ability to handle difficult samples.
   • Experiments show that AWM strategy is superior to random masking strategy.
-In addition, the noise suppression method proposed in this paper is applied to other models, and the performance is significantly improved.
-Among the AMNS methods, mAP index and mINP index are almost the best.
-Figure 1: An overview of our AMNS.
+In addition, the noise suppression method proposed in this paper is applied to other text-to-image frameworks, and the performance is significantly improved.
+Among the AMNS methods, mAP index and mINP index have achieved superior performance.
+Figure 1: Overview of our proposed framework AMNS.
 II METHOD
-The overall diagram of our proposed framework is shown in Fig.
-1, which consists of an image encoder ${\mathbf{E_{i}}}$, a text encoder ${\mathbf{E_{t}}}$ and an EMA-based image encoder $\overline{\mathbf{E_{i}}}$.
-Given an image, crop it and input it to $\overline{\mathbf{E_{i}}}$, a mask image is obtained using an AWM strategy to highlight image regions relevant to the textual description and fed into ${\mathbf{E_{i}}}$ together with the raw image, enabling ${\mathbf{E_{i}}}$ to extract a more accurate representation of the features.
-The model makes $\overline{\mathbf{E_{i}}}$ filter out features with lower weights by self-supervised learning, and combines with the proposed BSDM, which considers the inverse KL scatter, to inhibit the noisy labelling problem by inversely constraining the true label distribution through the predictive distribution.
-Meanwhile, combining with the proposed WAF further enhances the model’s attention and learning of hard samples by adjusting the weight coefficients.
-In addition, we adopt ID loss to aggregate feature representations of the same identity, which further improves the model retrieval performance.
+As shown in Fig.
+1, our framework’s overall structure includes image encoder ${\mathbf{E_{i}}}$, text encoder ${\mathbf{E_{t}}}$, and EMA-based image encoder $\overline{\mathbf{E_{i}}}$.
+The image is first cropped and processed by $\overline{\mathbf{E_{i}}}$ using AWM strategy to generate a mask that highlights regions relevant to the textual description.
+This masked image, combined with the original image, is then fed into ${\mathbf{E_{i}}}$ to extract more precise features.
+$\overline{\mathbf{E_{i}}}$ filters low-weight features via self-supervised learning, and with the proposed BSDM mitigates noisy label issues by aligning predictive and true distributions.
+Additionally, WAF enhances attention to hard samples by adjusting weight coefficients, ID loss aggregates the same identity features, further improving retrieval performance.
 Figure 2: BSDM diagram.
@@
 SDM loss[6] effectively differentiates matching from non-matching pairs by aligning the cosine similarity distribution $\mathbf{p}$ of image-text pairs with the ideal label distribution $\mathbf{q}$.
-However, due to noisy correspondences in the dataset, SDM forcibly fitting all positive sample pairs of a dataset causes the model to overfit data containing noise.
+However, the SDM loss overly focuses on negative samples while neglecting the learning of positive samples.
+Additionally, due to noisy correspondences in the dataset, forcibly fitting all positive sample pairs in the dataset leads to the model overfitting data containing noise.
 Therefore, we propose Bidirectional Similarity Distribution Matching (BSDM), the core idea of which is shown in Fig.
 2.
-Building upon the forward KL divergence, BSDM incorporates the inverse KL divergence, so that the distribution $\mathbf{p}$ can also exert constraints on the distribution $\mathbf{q}$.
-This bidirectional constraint helps alleviate Noisy Correspondences (NC) by allowing the model some flexibility in fitting the data, rather than blindly striving for perfect alignment with all positive pairs.
+BSDM allows the distribution p to also impose constraints on the distribution q.
+This bidirectional constraint provides the model with a certain degree of flexibility when fitting the data, enabling it to correctly fit positive sample pairs without blindly pursuing perfect alignment with all positive pairs, thereby helping to mitigate the impact of Noisy Correspondences (NC).
 0.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.873.073.073.073.073.573.573.573.574.074.074.074.074.574.574.574.575.075.075.075.0Mask RatioRank-1AMNSAWMAMNSRandom
-Figure 3: Variation of Rank-1 performance with different mask ratios for $m$ and $\tau$.
+Figure 3: Rank-1 performance changes of the two strategies at different mask ratios.
 First, given $N$ image-text pairs, $f_{i}^{v}$ and $f_{j}^{t}$ denote image global features and text global features.
@@
 $q_{i,j}=y_{i,j}/\sum_{k=1}^{N}y_{i,k}$ is the true matching probability.
-$KL(\mathbf{p_{i}}\parallel\mathbf{q_{i}})$ is the forward KL scatter, and the goal is to make $\mathbf{p_{i}}$ as close to $\mathbf{q_{i}}$ as possible, this helps the model to learn the correct data representation.
-$KL(\mathbf{q_{i}}\parallel\mathbf{p_{i}})$ is the inverse KL scatter.
+$KL(\mathbf{q_{i}}\parallel\mathbf{p_{i}})$ is the forward KL scatter, and the goal is to make $\mathbf{p_{i}}$ as close to $\mathbf{q_{i}}$ as possible, this helps the model to learn the correct data representation.
+$KL(\mathbf{p_{i}}\parallel\mathbf{q_{i}})$ is the inverse KL scatter.
 With the bidirectional KL scattering, the model is made to be pushed not only to adapt to the real labels that may contain noise, but also to allow the real labels to adapt to the model’s predictions, thus mitigating the effects of NC.
@@
 AMNSBSDM Best 74.25 89.54 93.83 67.67 53.16
-BSDM mitigates the effect of NC by introducing inverse KL scatter, but may weaken the model’s learning for hard samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples).
-Therefore, we consider the introduction of focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ we are able to adjust the model’s attention to the misclassified samples.
+BSDM combines forward KL and inverse KL to balance each other, thereby mitigating the impact of NC, but it may weaken the model’s ability to learn from hard samples (image-text similarity scores that are low, but are actually positive example samples, or similarity scores that are high, but are actually negative example samples).
+Therefore, we consider focal loss[25] and introduce proportional weight coefficients $\alpha$,$\beta$ in front of it to improve it, and by adjusting the coefficient factor $\gamma$ , we can adjust the model’s attention to the misclassified samples.
 We refer to this loss as Weight Adjustment Focal Loss (WAF).
@@
 $\mathcal{L}_{W\!AF}=\mathcal{L}_{W\!AF-i2t}+\mathcal{L}_{W\!AF-t2i}$ (7)
-By combining the two approaches of BSDM and WAF, the model not only suppresses the effect of noisy labels, but also enhances the learning of hard samples, thus achieving better performance in noisy data environments.
-Optimization.We also employ the commonly used ID loss[36], whose main purpose is to make feature representations of the same identity clustered together in the feature space, while feature representations of different identities are separated.
+By combining BSDM and WAF, the model suppresses noisy labels and enhances learning of hard samples, achieving better performance in noisy data environments.
+Optimization.We also employ the widely-used ID loss[36].
+This loss aims to cluster feature representations of the same identity within the feature space, while feature representations of different identities are separated.
 This helps the model to better distinguish between different entities, thus improving the performance of model retrieval.
 The overall optimization objective of our training is defined as:
-$\mathrm{L}=\mathrm{L}_{AWM}+\mathrm{L}_{BSDM}+\mathrm{L}_{W\!AF}+\mathrm{L}_{id}$ (8)
+$\mathcal{L}=\mathcal{L}_{AW\!M}+\mathcal{L}_{BSDM}+\mathcal{L}_{W\!AF}+\mathcal{L}_{id}$ (8)
 TABLE II: Performance comparison on three datasets. The best and second-best results are in bold and underline, respectively.
@@
 Table I shows that the use of BSDM loss in different methods can bring significant improvement in all aspects of performance, and the best performance is still achieved under the 20% and 50% noise rates, which fully indicates that the BSDM loss proposed by us is superior to the widely used TAL loss and SDM loss, and has better stability and robustness for NC.
-Table II shows that our method achieves good results in Rank metrics, and the mAP index and mINP index achieve the best results in the first and second datasets.
-Higher mINP metrics mean that the correctly matched images are indexed at a more advanced position, which fully validates that our method is better at finding all the correct matches and can be better used in real-life retrieval systems.
+Table II shows that our method achieves good results in Rank metrics.
+The mAP index and mINP index achieve the best results in the first and second datasets.
+Higher mINP metrics mean that the correctly matched images are indexed at a more advanced position, which fully validates that our method is better at finding the correct matches and can be better used in real-life retrieval systems.
 TABLE III: Ablation study on each component of AMNS on CUHK-PEDES.
@@
 In this section, we fine-tune the CLIP-ViT-B/16 model with AWM loss as a baseline and perform ablation studies on the CUHK-PEDES dataset.
-Effectiveness of BSDM: As can be seen from No.1 in Table III, just by adding BSDM to Baseline, the accuracies of Rank-1, Rank-5, and Rank-10 are improved by 4.22%, 2.21%, and 1.00%, respectively, and the values of mAP and mINP are improved by 3.08% and 2.68%, respectively, which is a good illustration of the BSDM’s Person Retrieval task in terms of its Effectiveness.
+Effectiveness of BSDM: As can be seen from No.1 in Table III, just by adding BSDM to Baseline, the accuracies of Rank-1, Rank-5, and Rank-10 are improved by 4.22%, 2.21%, and 1.00%, respectively, and the values of mAP and mINP are improved by 3.08% and 2.68%, respectively, which reflect the effectiveness of BSDM in the Person Retrieval task.
 Complementarity between BSDM and WAF: As can be seen from No.5 in Table III, the addition of WAF improves the Rank1 metric by 0.26%, and the mAP and mINP metrics by 0.31% and 0.45%, respectively, which fully demonstrates that WAF further improves the model’s ability to recognize hard samples when it is complemented with BSDM.
 IV Conclusion
-In this paper, we explore Noise Correspondence (NC) and the noisy matching problem caused by random masking strategy.
-We introduce Attention-Weighted Selective Mask to alleviate the noisy matching problem caused by random masking strategy, and propose a novel method, that is, Bidirectional Similarity Distribution Matching method, combined with the improved focal loss, effectively deal with the NC, and achieve excellent performance.
-We have carried out extensive experiments, which fully prove the superiority of the proposed method in solving the above problems, and prove that our method can find all the correct matches better.
+In this paper, we explore Noise Correspondence (NC) and the noisy matching problem.
+We introduce the Attention-Weighted Selective Mask (AWM) to alleviate the noise matching problem and propose a novel noise suppression method, Bidirectional Similarity Distribution Matching (BSDM), combined with an improved focal loss, to address the Noisy Correspondence (NC) issue.
+Together, these methods achieve excellent performance.
+Extensive experiments have been conducted, confirming the superiority of our approach in resolving the aforementioned issues and demonstrating its ability to find the correct matches better.